help I am stuck on this question
A financial institution offers personal loans with interest rates that vary from 8.25% to 12.5% compounded monthly, depending on a person's credit rating. Two people, one with the best credit rating and the other with the worst, take out loans of $12000 to be repaid in 30 monthly payments. How much interest does the person with the poor credit rating pay?

Answer by Boreal(15235)

formula Payment=PV*r/(1-(1+r))^-n, r rate, and n number compoundings
monthly it is PV(r/12)/(1-(1+r/12)^-nt), nt will be 30
=12000(0.125/12)/(1-(1+(.125/12)^-30))
=125/(1-(1.010416667)^-30) do without rounding until the end
$467.81 is the monthly payment
30 of them is $14034.30
the loan was $12000
The interest paid is the difference of $2034.30